OpenStudy (anonymous):

Solve the quadratic equation. 4x^2 - 30 = 0 A. ± √15 B. ± 15/√2 C. ± √15 /2 D. ± √30 /2

OpenStudy (anonymous):

\[4x^{2} - 30 = 0\]\[4x^{2} = 30\]\[x^{2}=\frac{ 30 }{ 4 }\]\[x=\pm \sqrt{\frac{ 30 }{ 4 }}=\pm \sqrt{\frac{ 15 }{ 2 }}=\pm \frac{ \sqrt{15} }{ \sqrt{2} }\]

OpenStudy (tkhunny):

\(\pm\dfrac{\sqrt{30}}{2}\)

OpenStudy (anonymous):

\[\pm \frac{ \sqrt{15} }{ \sqrt{2} }=\pm \frac{ \sqrt{30} }{ 2 }\] However, the latter is not one of the answer choices.

OpenStudy (tkhunny):

Except for D.

OpenStudy (anonymous):

Whoa, I must have been extremely tired yesterday. You are completely right.

OpenStudy (tkhunny):

No worries. The in-line notation is quite ambiguous. This is why I generally choose to code the LaTeX.